Calculate the concentration of C6H5NH2, H3O+, and OH− in a 0.230 M C6H5NH3Cl solution. (Kb(C6H5NH2)=3.9×10^−10


1
Expert's answer
2022-03-02T00:00:01-0500

Ka × kb = 10-14

ka = 10-14 ÷ 3.9×10-10

=2.564×10-5

Let [C6H5NH2] be x

[H3O+] is also x and

[C6H5NH3+] is (0.230-x)


Ka = ( [C6H5NH2] [H3O+] ÷ [C6H5NH3+] )

2.564×10-5 = x2 ÷ (0.230-x)

Solving for x quadratically,

X= 2.4251×10-3


Therefore, [H3O+] = 2.4251×10-3 M


[C6H5NH2] = 2.4251×10-3M


[C6H5NH3+] = 0.230 - 2.4251×10-3

= 0.2275M